# Joint Activity Detection and Channel Estimation for Massive Connectivity: Where Message Passing Meets Score-Based Generative Priors

### v1

@@ Abstract
Massive connectivity supports the sporadic access of a vast number of devices without requiring prior permission from the base station (BS). In such scenarios, the BS must perform joint activity detection and channel estimation (JADCE) prior to data reception. Message passing algorithms have emerged as a prominent solution for JADCE under a Bayesian inference framework. The existing message passing algorithms, however, typically rely on some hand-crafted and overly simplistic priors of the wireless channel, leading to significant channel estimation errors and reduced activity detection accuracy. In this paper, we focus on the problem of JADCE in a multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) grant-free random access network. We propose to incorporate a more accurate channel prior learned by score-based generative models into message passing, so as to push towards the performance limit of JADCE. Specifically, we develop a novel turbo message passing (TMP) framework that models the entire channel matrix as a super node, rather than factorizing it element-wise. This design enables the seamless integration of score-based generative models as a minimum mean-squared error (MMSE) denoiser. The variance of the denoiser, which is essential in message passing, can also be learned through score-based generative models. Our approach, termed score-based TMP for JADCE (STMP-JADCE), takes full advantages of the powerful generative prior and, meanwhile, benefits from the fast convergence speed of message passing. Numerical simulations show that STMP-JADCE drastically enhances the activity detection and channel estimation performance compared to the state-of-the-art baseline algorithms.
Index Terms:
Massive connectivity, joint activity detection and channel estimation (JADCE), score-based generative models, plug-and-play priors, message passing.
@@ I Introduction
Evolved from massive machine-type communication (mMTC) in fifth-generation (5G), massive communication (a.k.a. massive connectivity) has been identified as one of the six key usage scenarios in sixth-generation (6G) wireless networks [1]. A unique feature of massive communication is the sporadic transmission of short packets by massive machine-type devices, e.g., Internet-of-Things (IoT) devices, where only a small fraction of devices are active in packet transmission at any given time instant [2]. To support massive connectivity of machine-type devices, conventional grant-based random access schemes require each device to obtain a grant from the based station (BS) before transmission. The uplink grant request involves a complex handshaking process, which may result in significant signaling overhead and access latency [3, 4], especially in scenarios with massive connected devices.
@@
Due to the sporadic traffic of massive devices, activity detection and channel estimation can be formulated jointly as a compressive sensing (CS) problem. Message passing algorithms stand out as a powerful tool for efficient problem solving within a Bayesian framework. The performance of message passing relies heavily on the accuracy of the chosen priors, especially the prior of the wireless channel. Existing works on massive connectivity either assume a statistical channel model, such as Rayleigh fading [5, 6, 7, 8, 9] and Rician fading [10], or exploit channel sparsity in certain discrete transform domains, such as the discrete Fourier transform (DFT) angular basis [11, 12]. The former approach may be overly simplistic to model the multipath component, while the latter suffers from significant energy leakage caused by basis mismatch with the predefined codebook. The oversimplification or misspecification of the channel prior can result in substantial errors in channel estimation, which consequently degrades the accuracy of activity detection. Moreover, although not specifically tailored to massive connectivity, the authors in [13, 14] introduced advanced off-grid and grid-free line spectral estimation techniques for super-resolution channel estimation. Nevertheless, these methods continue to face challenges in resolving sub-paths within clustered multipath components and often entail prohibitive computational complexity.
@@
To speed up convergence, the authors in [24] proposed the integration of score-based generative models into message passing algorithms for solving inverse problems. Specifically, a score-based turbo message passing (STMP) framework was developed for compressive image recovery, in which the score network is integrated into Turbo-CS [25] as a plug-and-play denoiser. This framework inherits the rapid convergence of message passing, typically achieved within $10$ iterations, with each iteration requiring only a single-step denoising operation. Compared to DPS [18, 19, 20, 21], STMP eliminates the need for reverse diffusion sampling, and thus reduces the computational complexity by orders of magnitude. This motivates us to explore the feasibility of solving joint activity detection and channel estimation (JADCE) by integrating an accurate channel prior learned by score-based generative networks into message passing.
@@ I-A Contributions
In this paper, we consider a multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) grant-free random access network, where a massive number of devices sporadically access the uplink channel. We aim to develop an efficient message passing algorithm that incorporates a plug-and-play score-based generative channel prior to approach the performance limit of JADCE. The extension of STMP [24] for solving the JADCE problem is highly non-trivial, as the sporadic access of devices introduces additional activity-induced sparsity, which is not captured by the channel prior score network. To address this challenge, we propose a novel turbo message passing (TMP) framework that decouples the denoising of the channel and the device activity. This stands in contrast to prior works [5, 6, 11, 26], which treat the channel and its activity indicator as a whole. Moreover, unlike the separable denoisers in [5, 6, 11, 26] that operate element-wise for channel denoising, our TMP framework models the spatial-frequency channel matrix as a super node in the factor graph, enabling the incorporation of score-based generative models as non-separable denoisers for more effective exploitation of channel correlations. The score network enables single-step minimum mean-squared error (MMSE) denoising at each message passing iteration, supported by the well-established Tweedie’s formula [27, 28]. Additionally, we employ a second-order score network that outputs the Hessian diagonals of the log density to compute the variance of the denoiser. The plug-in implementation of the first- and second-order score networks yields the STMP algorithm tailored for JADCE, hence the name STMP-JADCE. The two score networks can be trained based on the denoising score matching technique [29, 15, 30]. The main contributions of this paper are summarized as follows.
@@
  • We develop the STMP-JADCE algorithm that seamlessly integrates score-based generative channel priors into the TMP framework. As a model-based deep learning solution, STMP-JADCE generalizes well across different device activity levels, pilot patterns, and signal-to-noise ratios (SNRs) without requiring retaining the score networks.
  • Experimental results show that our method significantly outperforms the existing message passing algorithms with a comparable or even faster convergence speed. Remarkably, our method consistently delivers over a $5$ dB normalized mean-squared error (NMSE) gain in channel estimation on the clustered delay line (CDL) channel models specified by the 3GPP standards [31].
@@ I-B Related Works
I-B1 Message Passing Algorithms for JADCE
It was shown in [5] that JADCE with single- and multi-antenna BSs corresponds to CS in single measurement vector (SMV) and multiple measurement vector (MMV) setups, respectively. The authors in [5] employed the approximate message passing (AMP) to solve both problems, in which the Rayleigh channel assumption is adopted to derive the MMSE denoiser in closed form. Interestingly, a follow-up work [6] revealed that the activity detection error by AMP is asymptotically zero as the number of BS antennas approaches infinity. Apart from [5, 6], the same Rayleigh channel model was also employed in [7, 8, 9] to facilitate message calculation and analysis. Considering that the line-of-sight (LoS) component in wave propagation is non-negligible in low Earth orbit (LEO) satellite networks, the authors in [10] proposed a Bernoulli-Rician message passing algorithm tailored for LEO satellite-enabled IoT applications with Rician channel modeling. The authors in [32] proposed a JADCE scheme exploiting the sparsity of the time-domain channel impulse response (CIR). However, an energy leakage problem arises from the inverse discrete Fourier transform (IDFT) to obtain the time-domain channel, which undermines its sparsity level. Considering frequency-domain channel estimation, the work in [11] developed a turbo-type algorithm that exploits the common angular sparsity over OFDM subcarriers. To reduce the pilot overhead, the authors in [26] constructed a block-wise linear model in the frequency domain, requiring the estimation of only the mean and slope in each sub-block. In the above work, the channel modeling and assigned prior distribution may fail to accurately capture the intricacies of the high-dimensional channel, leading to sub-optimal performance.
@@ I-B2 Model-Based and Model-Free Deep Learning for JADCE
It is well-known that message passing algorithms can be unfolded into a layer-wise structure analogous to a neural network (NN) [33]. By incorporating additional learnable parameters, the unfolded message passing networks can be trained using pairs of the channel samples, device activity realizations, and observations. It is shown that the deep unfolding technique can mitigate the potential message correlation issue on dense graphs and speed up convergence [8, 9, 12]. Nevertheless, deep unfolding, as a model-based deep learning approach, still requires prior knowledge of the wireless channel distribution. For example, the on-grid millimeter-wave parametric channel is adopted in [12], and the Rayleigh fading is adopted in [8, 9]. In contrast, pure data-driven approaches [34, 35] learn to fit certain input-output relationships without the need of explicit mathematical modeling. In [34], the authors proposed the use of two long short-term memory (LSTM) networks to learn the mappings from the received signals to the indices of active devices and to the channel coefficients. However, the existing deep learning-based solutions, whether model-based or model-free, are trained in a supervised manner using predefined active probabilities, specific pilot patterns, and fixed SNRs, which may not generalize effectively to diverse scenarios.
@@ I-C Organization and Notations
The remainder of this paper is organized as follows. Section II presents the system model, formulates the JADCE problem, and outlines the inherent limitations of traditional message passing algorithms for problem-solving. Section III develops a TMP framework that supports the integration of score-based generative priors. Section IV reveals the relationship between the score function and MMSE denoising, and then specifies the training objectives for the score networks. In Section V, we present the STMP-JADCE algorithm and related engineering tricks. In Section VI, we provide extensive simulation results to validate the efficiency and effectiveness of the proposed method. Finally, we conclude this paper in Section VII.
@@ II-A System Model
where $\mathbf{Q}\triangleq\left[\mathbf{Q}_{1},\dots,\mathbf{Q}_{K}\right]\in\mathbb{C}^{NT\times NK}$, and $\mathbf{X}\triangleq\left[\mathbf{X}_{1}^{\sf H},\dots,\mathbf{X}_{K}^{\sf H}\right]^{\sf H}\in\mathbb{C}^{NK\times M}$. For simplicity, we impose the condition that $\mathbf{Q}$ is partial orthogonal, specifically requiring that $\mathbf{Q}\mathbf{Q}^{\sf H}=KP\mathbf{I}_{NT}$, where each device has the same transmit power budget $P$. The design of partial orthogonal $\mathbf{Q}$ is deferred to Section V-C. Due to the sporadic access of the devices, matrix $\mathbf{X}$ exhibits a block-sparsity structure. Specifically, if the $k$-th device is inactive, we have $\mathbf{X}_{k}=\alpha_{k}\mathbf{H}_{k}=\mathbf{0}$.
@@ II-B Problem Statement
Alternatively, message passing algorithms [6, 5, 26, 46] provide a low-complexity iterative approach for approximately calculating the MMSE estimator. In message passing, we factorize the joint posterior distribution according to Baye’s theorem and the conditional independence relationships among the variables, expressed as
@@
where $\mathbf{x}_{m}\in\mathbb{C}^{NK}$ and $\mathbf{y}_{m}\in\mathbb{C}^{NT}$ denote the $m$-th column of $\mathbf{X}=\left[\mathbf{x}_{1},\dots,\mathbf{x}_{M}\right]$ and $\mathbf{Y}=\left[\mathbf{y}_{1},\dots,\mathbf{y}_{M}\right]$, respectively. A factor graph is then utilized to represent the joint distribution in (5), as illustrated in Fig. 2. The main difficulty in implementing message passing on the factor graph lies in the message calculation associated with the factor node $p(\mathbf{H}_{k})$, detailed as follows.
@@
To summarize, the message passing framework faces a dilemma: the adopted channel prior is either too imprecise to ensure good performance, or too complex to facilitate efficient message calculation. This paper seeks to address the aforementioned challenges by plugging in deep generative priors for efficient message calculation. We begin by developing a TMP framework for JADCE. Then, we present a unique interpretation of the belief calculation at factor node $p(\mathbf{H}_{k})$ as an MMSE denoising problem under AWGN observations. This insight enables the integration of score-based generative networks for provable MMSE denoising, eliminating the need to compute high-dimensional integrals.
TABLE I: Factor Nodes and Their Abbreviations

### v2

@@ Abstract
Massive connectivity supports the sporadic access of a vast number of devices without requiring prior permission from the base station (BS). In such scenarios, the BS must perform joint activity detection and channel estimation (JADCE) prior to data reception. Message-passing algorithms have emerged as a prominent solution for JADCE under a Bayesian inference framework. The existing message-passing algorithms, however, typically rely on some hand-crafted and overly simplistic priors of the wireless channel, leading to significant channel estimation errors and reduced activity detection accuracy. In this paper, we focus on the problem of JADCE in a multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) grant-free random access network. We propose to incorporate a more accurate channel prior learned by score-based generative models (a.k.a. diffusion models) into message passing, so as to push towards the performance limit of JADCE. Specifically, we develop a novel turbo message passing (TMP) framework that models the entire channel matrix as a super node, rather than factorizing it element-wise. This design enables the seamless integration of score-based generative models as a minimum mean-squared error (MMSE) denoiser. The variance of the denoiser, which is essential in message passing, can also be learned through score-based generative models. Our approach, termed score-based TMP for JADCE (STMP-JADCE), takes full advantages of the powerful generative prior and, meanwhile, benefits from the fast convergence speed of message passing. Numerical simulations show that STMP-JADCE drastically enhances the activity detection and channel estimation performance compared to the state-of-the-art baseline algorithms.
@@ I Introduction
Evolved from massive machine-type communication (mMTC) in fifth-generation (5G), massive communication (a.k.a. massive connectivity) has been identified as one of the six key usage scenarios in sixth-generation (6G) wireless networks [1]. A unique feature of massive communication is the sporadic transmission of short packets by massive machine-type devices, e.g., Internet-of-Things (IoT) devices, where only a small fraction of devices are active in packet transmission at any given time instant [2]. To support massive connectivity of machine-type devices, conventional grant-based random access schemes require each device to obtain a grant from the base station (BS) before transmission. The uplink grant request involves a complex handshaking process, which may result in significant signaling overhead and access latency [3, 4], especially in scenarios with massive connected devices.
@@
Due to the sporadic traffic of massive devices, activity detection and channel estimation can be formulated jointly as a compressive sensing (CS) problem. Message-passing algorithms stand out as a powerful tool for efficient problem solving within a Bayesian framework. The performance of message passing relies heavily on the accuracy of the chosen priors, especially the prior of the wireless channel. Existing works on massive connectivity either assume a statistical channel model, such as Rayleigh fading [5, 6, 7, 8, 9] and Rician fading [10], or exploit channel sparsity in certain discrete transform domains, such as the discrete Fourier transform (DFT) angular basis [11, 12]. The former approach may be overly simplistic to model the multipath component, while the latter suffers from significant energy leakage caused by basis mismatch with the predefined codebook. The oversimplification or misspecification of the channel prior can result in substantial errors in channel estimation, which consequently degrades the accuracy of activity detection. Moreover, although not specifically tailored to massive connectivity, the authors in [13, 14] introduced advanced off-grid and grid-free line spectral estimation techniques for super-resolution channel estimation. Nevertheless, these methods continue to face challenges in resolving sub-paths within clustered multipath components and often entail prohibitive computational complexity.
@@
To speed up convergence, the authors in [24, 25] proposed the integration of score-based generative models into message-passing algorithms for solving inverse problems. Specifically, a score-based turbo message passing (STMP) framework was developed for compressive image recovery, in which the score network is integrated into Turbo-CS [26] as a plug-and-play denoiser. This framework inherits the rapid convergence of message passing, typically achieved within $10$ iterations, with each iteration requiring only a single-step denoising operation. Compared to DPS [18, 19, 20, 21, 22, 23], STMP eliminates the need for reverse diffusion sampling, and thus reduces the computational complexity by orders of magnitude. This motivates us to explore the feasibility of solving joint activity detection and channel estimation (JADCE) by integrating an accurate channel prior learned by score-based generative networks into message passing.
@@ I-A Contributions
In this paper, we consider a multiple-input multiple-output orthogonal frequency division multiplexing (MIMO-OFDM) grant-free random access network, where a massive number of devices sporadically access the uplink channel. We aim to develop an efficient message-passing algorithm that incorporates a plug-and-play score-based generative channel prior to approach the performance limit of JADCE. The extension of STMP [24, 25] for solving the JADCE problem is highly non-trivial, as the sporadic access of devices introduces additional activity-induced sparsity, which is not captured by the channel prior score network. To address this challenge, we propose a novel turbo message passing (TMP) framework that decouples the denoising of the channel and the device activity. This stands in contrast to prior works [5, 6, 11, 27], which treat the channel and its activity indicator as a whole. Moreover, unlike the separable denoisers in [5, 6, 11, 27] that operate element-wise for channel denoising, our TMP framework models the spatial-frequency channel matrix as a super node in the factor graph, enabling the incorporation of score-based generative models as non-separable denoisers for more effective exploitation of channel correlations. The score network enables single-step minimum mean-squared error (MMSE) denoising at each message-passing iteration, supported by the well-established Tweedie’s formula [28, 29]. Additionally, we employ a second-order score network that outputs the Hessian diagonals of the log density to compute the variance of the denoiser. The plug-in implementation of the first- and second-order score networks yields the STMP algorithm tailored for JADCE, hence the name STMP-JADCE. The two score networks can be trained based on the denoising score matching technique [30, 15, 31]. The main contributions of this paper are summarized as follows.
@@
  • We develop the STMP-JADCE algorithm that seamlessly integrates score-based generative channel priors into the TMP framework. As a model-based deep learning solution, STMP-JADCE generalizes well across different device activity levels, pilot patterns, and signal-to-noise ratios (SNRs) without requiring retraining the score networks.
  • Experimental results show that our method significantly outperforms the existing message-passing algorithms with a comparable or even faster convergence speed. Remarkably, our method consistently delivers over a $5$ dB normalized mean-squared error (NMSE) gain in channel estimation on the clustered delay line (CDL) channel models specified by the 3GPP standards [32].
Beyond the algorithmic advances, this work underscores a conceptual shift in Bayesian inference for massive connectivity: from traditional model-based methods relying on hand-crafted statistical priors (e.g., Rayleigh or Rician fading), to data-driven approaches that directly learn channel distributions from data via score-based generative modeling. This philosophical transition is central to our contribution, as it enables Bayesian inference to harness expressive learned priors without sacrificing the efficiency of message passing.
@@ I-B Related Works
I-B1 Message-Passing Algorithms for JADCE
It was shown in [5] that JADCE with single- and multi-antenna BSs corresponds to CS in single measurement vector (SMV) and multiple measurement vector (MMV) setups, respectively. The authors in [5] employed the approximate message passing (AMP) to solve both problems, in which the Rayleigh channel assumption is adopted to derive the MMSE denoiser in closed form. Interestingly, a follow-up work [6] revealed that the activity detection error by AMP is asymptotically zero as the number of BS antennas approaches infinity. Apart from [5, 6], the same Rayleigh channel model was also employed in [7, 8, 9] to facilitate message calculation and analysis. Considering that the line-of-sight (LoS) component in wave propagation is non-negligible in low Earth orbit (LEO) satellite networks, the authors in [10] proposed a Bernoulli-Rician message-passing algorithm tailored for LEO satellite-enabled IoT applications with Rician channel modeling. The authors in [33] proposed a JADCE scheme exploiting the sparsity of the time-domain channel impulse response (CIR). However, an energy leakage problem arises from the inverse discrete Fourier transform (IDFT) to obtain the time-domain channel, which undermines its sparsity level. Considering frequency-domain channel estimation, the work in [11] developed a turbo-type algorithm that exploits the common angular sparsity over OFDM subcarriers. To reduce the pilot overhead, the authors in [27] constructed a block-wise linear model in the frequency domain, requiring the estimation of only the mean and slope in each sub-block. In the above work, the channel modeling and assigned prior distribution may fail to accurately capture the intricacies of the high-dimensional channel, leading to sub-optimal performance.
@@ I-B2 Model-Based and Model-Free Deep Learning for JADCE
It is well-known that message-passing algorithms can be unfolded into a layer-wise structure analogous to a neural network (NN) [34]. By incorporating additional learnable parameters, the unfolded message passing networks can be trained using pairs of the channel samples, device activity realizations, and observations. It is shown that the deep unfolding technique can mitigate the potential message correlation issue on dense graphs and speed up convergence [8, 9, 12]. Nevertheless, deep unfolding, as a model-based deep learning approach, still requires prior knowledge of the wireless channel distribution. For example, the on-grid millimeter-wave parametric channel is adopted in [12], and the Rayleigh fading is adopted in [8, 9]. In contrast, pure data-driven approaches [35, 36] learn to fit certain input-output relationships without the need of explicit mathematical modeling. In [35], the authors proposed the use of two long short-term memory (LSTM) networks to learn the mappings from the received signals to the indices of active devices and to the channel coefficients. However, the existing deep learning-based solutions, whether model-based or model-free, are trained in a supervised manner using predefined active probabilities, specific pilot patterns, and fixed SNRs, which may not generalize effectively to diverse scenarios.
@@ I-C Organization and Notations
The remainder of this paper is organized as follows. Section II presents the system model, formulates the JADCE problem, and outlines the inherent limitations of traditional message-passing algorithms for problem-solving. Section III develops a TMP framework that supports the integration of score-based generative priors. Section IV reveals the relationship between the score function and MMSE denoising, and then specifies the training objectives for the score networks. In Section V, we present the STMP-JADCE algorithm and related engineering tricks. In Section VI, we provide extensive simulation results to validate the efficiency and effectiveness of the proposed method. Finally, we conclude this paper in Section VII.
@@ II-A System Model
where $\mathbf{Q}\triangleq\left[\mathbf{Q}_{1},\dots,\mathbf{Q}_{K}\right]\in\mathbb{C}^{NT\times NK}$, and $\mathbf{X}\triangleq\left[\mathbf{X}_{1}^{\sf H},\dots,\mathbf{X}_{K}^{\sf H}\right]^{\sf H}\in\mathbb{C}^{NK\times M}$. For simplicity, we impose the condition that $\mathbf{Q}$ is partial orthogonal, specifically requiring that $\mathbf{Q}\mathbf{Q}^{\sf H}=KP\mathbf{I}_{NT}$, where each device has the same transmit power budget $P$. The design of partial orthogonal $\mathbf{Q}$ is deferred to Section V-C. Due to the sporadic access of the devices, matrix $\mathbf{X}$ exhibits a block-sparsity structure. Specifically, if the $k$-th device is inactive, we have $\mathbf{X}_{k}=\alpha_{k}\mathbf{H}_{k}=\mathbf{0}$. Frequently used notations in this paper are summarized in Table I.
TABLE I: Summary of Notations
Notation Description
$M$ Number of BS antennas
NN Number of OFDM subcarriers
$K$ Number of sporadic access devices
$T$ Number of OFDM symbols
$P$ Transmit power budget of each device
$\lambda$ Prior probability that a device is active
$\alpha_{k}$ Indicator of whether device $k$ is active
$\mathbf{H}_{k}$ Channel matrix from device $k$ to the BS
$\mathbf{X}_{k}$ Effective channel matrix $\mathbf{X}_{k}\triangleq\alpha_{k}\mathbf{H}_{k}$
$\mathbf{h}_{km}$ ($\mathbf{x}_{km}$) The $m$-th column of $\mathbf{H}_{k}$ ($\mathbf{X}_{k}$)
$\mathbf{X}$ $\mathbf{X}\triangleq\left[\mathbf{X}_{1}^{\sf H},\dots,\mathbf{X}_{K}^{\sf H}\right]^{\sf H}$
$\mathbf{x}_{m}$ The $m$-th column of $\mathbf{X}$
$\mathbf{Q}$ Concatenated pilot symbol matrix
$\mathbf{Y}$ Concatenated received signal matrix
$\mathbf{N}$ Concatenated AWGN matrix
$\mathbf{x}_{m,\mathsf{A}}^{\mathtt{pri}}$, $v_{m,\mathsf{A}}^{\mathtt{pri}}$ Prior mean and variance of $\mathbf{x}_{m}$ in module A
$\mathbf{x}_{m,\mathsf{A}}^{\mathtt{post}}$, $v_{m,\mathsf{A}}^{\mathtt{post}}$ Posterior mean and variance of $\mathbf{x}_{m}$ in module A
$\mathbf{x}_{m,\mathsf{A}}^{\mathtt{ext}}$, $v_{m,\mathsf{A}}^{\mathtt{ext}}$ Extrinsic mean and variance of $\mathbf{x}_{m}$ in module A
$\mathbf{x}_{km,\mathsf{B}}^{\mathtt{pri}}$, $v_{m,\mathsf{B}}^{\mathtt{pri}}$ Prior mean and variance of $\mathbf{x}_{km}$ in module B
$\mathbf{x}_{km,\mathsf{B}}^{\mathtt{post}}$, $v_{m,\mathsf{B}}^{\mathtt{post}}$ Posterior mean and variance of $\mathbf{x}_{km}$ in module B
$\mathbf{x}_{km,\mathsf{B}}^{\mathtt{ext}}$, $v_{m,\mathsf{B}}^{\mathtt{ext}}$ Extrinsic mean and variance of $\mathbf{x}_{km}$ in module B
$\mathbf{h}_{km}^{\mathtt{pri}}$, $\tau_{m}^{\mathtt{pri}}$ Prior mean and variance of $\mathbf{h}_{km}$ in module B
$\mathbf{h}_{km}^{\mathtt{post}}$, $\tau_{m}^{\mathtt{post}}$ Posterior mean and variance of $\mathbf{h}_{km}$ in module B
$\mathbf{h}_{km}^{\mathtt{ext}}$, $\tau_{m}^{\mathtt{ext}}$ Extrinsic mean and variance of $\mathbf{h}_{km}$ in module B
$\lambda_{k}^{\mathtt{post}}$ Posterior estimate of the active probability of device $k$
@@ II-B Problem Statement
Alternatively, message-passing algorithms [6, 5, 27, 47] provide a low-complexity iterative approach for approximately calculating the MMSE estimator. In message passing, we factorize the joint posterior distribution according to Baye’s theorem and the conditional independence relationships among the variables, expressed as
@@
where $\mathbf{x}_{m}\in\mathbb{C}^{NK}$ and $\mathbf{y}_{m}\in\mathbb{C}^{NT}$ denote the $m$-th column of $\mathbf{X}=\left[\mathbf{x}_{1},\dots,\mathbf{x}_{M}\right]$ and $\mathbf{Y}=\left[\mathbf{y}_{1},\dots,\mathbf{y}_{M}\right]$, respectively. A factor graph is then utilized to represent the joint distribution in (5), as illustrated in Fig. 2. The distributions of each factor node and their abbreviations are summarized in Table II. The main difficulty in implementing message passing on the factor graph lies in the message calculation associated with the factor node $p(\mathbf{H}_{k})$, detailed as follows.
@@
To summarize, the message-passing framework faces a dilemma: the adopted channel prior is either too imprecise to ensure good performance, or too complex to facilitate efficient message calculation. This paper seeks to address the aforementioned challenges by plugging in deep generative priors for efficient message calculation. We begin by developing a TMP framework for JADCE. Then, we present a unique interpretation of the belief calculation at the factor node $p(\mathbf{H}_{k})$ as an MMSE denoising problem under AWGN observations. This insight enables the integration of score-based generative networks for provable MMSE denoising, eliminating the need to compute high-dimensional integrals.
TABLE II: Factor Nodes and Their Abbreviations